# Influence of initiators on the tipping point in the extended Watts model

### v1

@@ IV Summary
In this study, we investigated the impact of initiators (seeds) on the tipping point in the cascade dynamics of the Watts model in networks. To this end, we employed the extended Watts model, where seed neighbors (nodes with connections to initiators) have a different adoption threshold than ordinary nodes (nodes without connections to initiators). By applying the tree approximation to the extended Watts model on ERRGs, we determined the cascade sizes for various combinations of adoption thresholds for seed neighbors and ordinary nodes. The results indicate that the tipping point, representing a sudden increase in the active node fraction in the final state, depends on two factors: whether a giant SN-component is formed when the seed fraction is large enough to trigger cascades among seed neighbors, and whether the giant ON-component is maintained when newly activated nodes trigger further activations among ordinary nodes. We derived the size of the giant component formed by each type of node and its associated critical point, demonstrating how giant components appear or disappear with increasing seed fraction. The coexistence of a giant SN-component and giant ON-component suggests that large cascades on each giant component can occur at different tipping points. We presented an example in which a tipping point appears twice when there is a gap between the adoption thresholds of seed neighbors and ordinary nodes. Through the size distribution of connected components of active nodes of different types, we observed that the initial cascade at a smaller tipping point occurs exclusively in the giant SN-component, while the subsequent cascade at a larger tipping point occurs in the giant ON-component as well.

### v2

@@ IV Summary
Figure 7: Critical point $\rho_{0,{\rm c}}^{\rm S}$ for S-components (black dot-dashed line), $\rho_{0,{\rm c}}^{\rm SN}$ for SN-components (red dotted line), $\rho_{0,{\rm c}}^{\rm SN+ON}$ for (SN+ON)-components (green solid line), and $\rho_{0,{\rm c}}^{\rm ON}$ for ON-components (blue dashed line) in the $(\rho_{0},\gamma)$ plane for random scale-free networks with $p_{k}=k^{-\gamma}/\sum_{k^{\prime}=k_{\rm min}}^{\infty}k^{\prime-\gamma}$. Here we set $k_{\rm min}=2$. Two critical points for SN-components exist when $\gamma<\gamma_{c}\approx 3.825$.
In this study, we investigated the impact of initiators (seeds) on the tipping point in the cascade dynamics of the Watts model in networks. To this end, we employed the extended Watts model, where seed neighbors (nodes with connections to initiators) have a different adoption threshold than ordinary nodes (nodes without connections to initiators). By applying the tree approximation to the extended Watts model on ERRGs, we determined the cascade sizes for various combinations of adoption thresholds for seed neighbors and ordinary nodes. The results indicate that the tipping point, representing a sudden increase in the active node fraction in the final state, depends on two factors: whether a giant SN-component is formed when the seed fraction is large enough to trigger cascades among seed neighbors, and whether the giant ON-component is maintained when newly activated nodes trigger further activations among ordinary nodes. We derived the size of the giant component formed by each type of node and its associated critical point, demonstrating how giant components appear or disappear with increasing seed fraction. When the average degree of the ERRG is large ($\langle k\rangle>\langle k\rangle_{c}\approx 2.580$), both the giant SN-component and the giant ON-component exist within a certain range of $\rho_{0}$. The coexistence of a giant SN-component and giant ON-component suggests that large cascades on each giant component can occur at different tipping points. We presented an example in which a tipping point appears twice when there is a gap between the adoption thresholds of seed neighbors and ordinary nodes. Through the size distribution of connected components of active nodes of different types, we observed that the initial cascade at a smaller tipping point occurs exclusively in the giant SN-component, while the subsequent cascade at a larger tipping point occurs in the giant ON-component as well.
In this study, we focused solely on ERRGs. An interesting question is whether the results presented here hold for other network topologies, particularly for scale-free networks. The behavior of the extended Watts model on scale-free networks warrants further investigation due to their degree heterogeneity and the significant differences between nodes constituting the SN-components and ON-components under random seed placement. We note that the formation of each type of giant component is strongly influenced by the degree heterogeneity. Figure 7 plots the critical points for each type of giant component in the $(\rho_{0},\gamma)$ plane for random scale-free networks with $p_{k}=k^{-\gamma}/\sum_{k^{\prime}=k_{\rm min}}^{\infty}k^{\prime-\gamma}$, obtained using the generating function approach described in Appendix A. When seeds are placed randomly in scale-free networks, seed neighbors are more likely to include high-degree nodes, while ordinary nodes are less likely to do so. Consequently, the giant SN-component easily forms for $\gamma<\gamma_{c}\approx 3.825$, whereas the giant ON-component disintegrates with an increasing seed fraction. It remains unclear and worth further discussion whether a combination of adoption thresholds exists that could result in the emergence of two tipping points in scale-free networks.
@@ Acknowledgment
Appendix B Case where $\theta_{\rm SN}>\theta_{\rm ON}$
Figure 8: Active node fraction $\rho_{\infty}$ evaluated from Eq. (4) and Eq. (5) for the extended Watts model on ERRGs with $\langle k\rangle=10$. The black dashed line, red dotted line, and blue solid line represent the results when $(\theta_{\rm SN},\theta_{\rm ON})=(0.5,0.2)$, $(\theta_{\rm SN},\theta_{\rm ON})=(0.6,0.1)$, and $(\theta_{\rm SN},\theta_{\rm ON})=(0.7,0.0)$, respectively. The symbols represent the simulation results for ERRGs with $10^{6}$ nodes, averaged over $10^{2}$ samples ($10$ graphs $\times$ $10$ initial states). The theoretical lines are in good agreement with the simulation results.
In the main text, we concentrated on the case where $\theta_{\rm SN}<\theta_{\rm ON}$. In this appendix, we address the case where $\theta_{\rm SN}>\theta_{\rm ON}$, indicating that ordinary nodes are more easily activated than seed neighbors. We consider the extended Watts model on ERRGs with $\theta_{\rm SN}>\theta_{\rm ON}$. If activations propagate among seed neighbors, ordinary nodes will also rapidly join the cascade since they are more easily activated than seed neighbors. An interesting scenario arises when $\theta_{\rm SN}$ is significantly high and $\theta_{\rm ON}$ is considerably low. While a significant number of seeds is required for activations to propagate among seed neighbors having a high adoption threshold, a large cascade that spreads across the entire network is more likely to occur when seed fraction $\rho_{0}$ is small.
Figure 8 shows the active node fraction $\rho_{\infty}$ as a function of seed fraction $\rho_{0}$ for the extended Watts model on an ERRG with $\langle k\rangle=10$, plotted for three different combinations of $\theta_{\rm SN}$ and $\theta_{\rm ON}(<\theta_{\rm SN})$. In all cases, the theoretical lines obtained from the tree approximation (Eq. (4) and Eq. (5)) agree well with the results of the Monte Carlo simulations. The black dashed line for $\theta_{\rm SN}=0.5$ and $\theta_{\rm ON}=0.2$ ($\theta_{\rm SN}$ is not too high and $\theta_{\rm ON}$ is not too low) indicates that a large cascade, where $\rho_{\infty}$ jumps from $\rho_{\infty}\approx\rho_{0}$ to $\rho_{\infty}=1$, occurs at a certain tipping point $\rho_{0}\approx 0.24$ with increasing $\rho_{0}$.
The interesting case where $\theta_{\rm SN}=0.6$ and $\theta_{\rm ON}=0.1$ is illustrated by the red dotted line in Fig. 8. We find that $\rho_{\infty}$ behaves non-monotonically with an increasing seed fraction $\rho_{0}$: $\rho_{\infty}=1$ when $\rho_{0}$ is small; as $\rho_{0}$ increases, $\rho_{\infty}$ drops discontinuously at a certain seed fraction; $\rho_{\infty}$ decreases and falls to $\rho_{\infty}\approx\rho_{0}$ as $\rho_{0}$ continues to increase; and with further increases in $\rho_{0}$, $\rho_{\infty}$ increases again and jumps to $\rho_{\infty}=1$ at a tipping point.
When $\rho_{0}$ is small, most nodes are members of the giant ON-component. Even when the activation threshold for seed neighbors is very high, a large cascade that spreads across the entire network is possible if random seed placement leads to a configuration where even a very small number of seed neighbors are surrounded by enough seeds to trigger activation (Monte Carlo simulations sample such a configuration each time when networks are sufficiently large). If $\theta_{\rm ON}$ is sufficiently low, even the activation of a single seed neighbor can trigger a large cascade within the giant ON-component, leading to the further activation of seed neighbors by activated ordinary nodes and resulting in all nodes becoming active.
As the seed fraction increases, the active node fraction drops discontinuously at a certain seed fraction ($\rho_{0}\approx 0.13$). This reflects the formation of the giant SN-component and the increased connectivity between seed neighbors, preventing the activation of ordinary nodes from spreading to the giant SN-component. It should be noted that the size of this drop is approximately equal to the fraction of the giant SN-component. As the seed fraction continues to increase, the active node fraction gradually decreases, reaching a minimum at a certain seed fraction ($\rho_{0}\approx 0.16$). This roughly coincides with the disappearance of the giant ON-component. In the absence of the giant ON-component, even if some seed neighbors are activated by chance in random seed placement, the propagation of activation to ordinary nodes is negligibly limited. Subsequently, the active node fraction increases continuously with the seed fraction, and jumps to 1 at a tipping point ($\rho_{0}\approx 0.39$), indicating the onset of a large cascade on the giant SN-component.
The case where $\theta_{\rm SN}=0.7$ and $\theta_{\rm ON}=0.0$ (the blue solid line) is even clearer. In this setting, all ordinary nodes are always active due to their zero adoption threshold. From the figure, we find again that a large cascade which spreads across the entire network occurs for small $\rho_{0}$ and the active node fraction drops discontinuously at a larger seed fraction ($\rho_{0}\approx 0.07$). For the latter, a plausible explanation is that the connection between seed neighbors in the giant SN-component is strengthened, preventing seed neighbors within it from being activated by ordinary nodes (which all become active) and seeds.